A sample of nitrogen occupies 11.2 liters under a pressure of 580 torr at 32◦C. What volume would it occupy at 32◦C if the pressure were increased to 810 torr?


1
Expert's answer
2021-03-09T06:03:27-0500

pV=nRT.

n=pV/RT = 77.327 x 11.2/8.314 x 305 =

0,34153823 moles.

V=nRT/p= 0,34153823 x 8.314 x 305/107.991 = 8,01976457 L.